Suppose X has a continuous uniform distribution over the interval Round your answers to 3 decimal places. (a) Determine the mean, variance, and standard deviation of X 1-1, 」 Mean- 0.33 Variance- Standard deviation-(0.574 (b) Determine the value for x such that 0.82
Answer #1

Solution :

\mathrm{Let\:\:X\:\:be\:a\:random\:variable\:with\:continuous\:uniform\:distribution\:over\:the\:interval\:\left[-1,\:1\right]\:with\:the\:pdf\::}

f(x)=\frac{1}{1-(-1)}=\frac{1}{2},\:\:-1<x<1

( a )

\mathrm{Mean\:\:E(X)}=\frac{-1+1}{2}=0

\mathrm{Var(X)}=\frac{(1-(-1))^{2}}{12}=\frac{(2)^{2}}{12}=\frac{4}{12}=\frac{1}{3}=0.33333\approx 0.333

\mathrm{Std(X)}=\sqrt{\left ( \frac{1}{3} \right )}=0.57735\approx 0.577

----------------------------------------------------------------------

( b )

\mathrm{P}(-x<\mathrm{X}<x)=\int_{-x}^{x}\frac{1}{2}dx=\frac{1}{2}*2x=x=0.82